# Periodic Chains Scheduling on Dedicated Resources - A Crucial Problem in Time-Sensitive Networks

### v1

@@ 4.3 Constraint Programming Matheuristic
$ps_{i}^{k},pe_{i}^{k},df_{i}^{k}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathtt{integerVar}$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k}\in\mathcal{C}^{k}$ (21)
@@
Eq. 9 ensures the non-overlapping of intervals on each resource. Eqs. 10, 11 and 12 either fix or limit the intervals of the tasks. Variables $ps_{i}^{k}$ and $pe_{i}^{k}$ correspond to the start- and completion time of the task once the resource offset is considered. We used them to calculate the consecutive differences within the chain $df_{i}^{k}$. This, together with tasks’ processing times, forms the start-to-end latency $S^{k}$ in equation Eq. 18.
@@
Criterion precisely models the degeneracy $D_{sum}$. Note that both $D_{\max},D_{sum}({\alpha})$ can be modeled in a similar manner. Furthermore, we add the following constraint for sum-based criteria to ensure that the degeneracy of individual chains is not worsened, which we observed to improve the performance. Let $deg^{k}$ be the degeneracy of chain $\mathcal{C}^{k}$ given the current solution. Then we add the following constraints:
$S^{k}\leq(deg^{k}+1)\cdot\tau(\beta)$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}$ (22)
@@
For large instances, the performance of the CP solver starts to diminish due to the large number of task intervals. Thus, when the size of the problem exceeds a defined bound, we do not optimize all the chains of given period $\tau(\beta)$ simultaneously; rather, we optimize only $K$ chains with the largest value of degeneracy (keeping others fixed). We repeat this with $K$ most problematic chains until a time limit is reached.
@@ 5 Experiments
The experiments were primarily performed on synthetically generated instances inspired by those in [30]. Initially, each resource is split into segments that become tasks of period $\tau(1)$ and segments that are recursively split further to allocate longer-period tasks. Base period $\tau(1)$ was set to 100, 200, or 400, and there were between 3 to 5 distinct periods per instance, with ratios of 2, 3, and 4 between consecutive periods. Chains are generated afterward, ensuring that the degeneracy of the instance is zero.
@@
Figure 8: Topologies used for TOP instance sets.
Three kinds of instance sets were produced. Set GEN-x contains the instances for the general case, where the chain can pass from one arbitrary resource to another, even returning to the original one; processing times among tasks of the chain can differ, and x refers to $\min_{m}U_{m}$; the minimum utilization among the resources.
Sets TOP were generated using underlying topologies: star, triangle, bridge, and stubbed line (Fig. 8). TOP-L contains instances using the stubbed line topology; the rest is put into the set TOP-R. Processing time is the same for all tasks of each chain. The links are full-duplex, and messages are sent between endpoints. However, there are also tasks being sent between two neighboring nodes only to increase the utilization (when we omit these additional tasks in sets TOP-R, TOP-L, we obtain less-utilized sets TOP-R-L, TOP-L-L).
Finally, THEORY set is used to discuss our complexity results; it consists of line and out-tree topologies, where all chains start in the same common resource (see Section 3.2). The parameters of the sets are presented in Table 1.
# instances # resources avg # tasks avg # chains avg chain length avg $\min_{m}U_{m}$ GEN-1 1194 5-10 2193 254 8.4 1.00 GEN-0.98 1194 5-10 1786 208 8.3 0.98 GEN-0.96 1194 5-10 1558 183 8.2 0.96 GEN-0.94 1194 5-10 1374 163 8.1 0.94 GEN-0.9 1194 5-10 1093 131 7.9 0.90 TOP-L 796 46 7634 2247 2.7 1.00 TOP-L-L 796 46 6155 2246 2.7 0.71 TOP-R 2388 10-24 2487 792 2.3 1.00 TOP-R-L 2388 10-24 1929 792 2.35 0.73 THEORY 320 10-77 1102 96 11.2 0.66
@@ 5.1 Finding Feasible Solutions
First, we studied the effect of a warm start and the ability of the local search to find feasible and high-quality solutions. We ended with a strategy that uses the CP to warm-start the heuristic for instances with utilization close to 100%; otherwise, it relies on the local search itself. This strategy provided a good trade-off between the feasibility success rate and the value of the criterion.
The experiment was performed on TOP and GEN instances with a time limit of 6 minutes. In Table 2, we report the percentage of instances where the feasible solution was found, the average value of degeneracy $D_{sum}$, and a percentage of instances with 0-$D_{sum}$ solutions (when the final solution has degeneracy equal to 0). The results are reported for both leftmost and predecessor heuristic methods, either being warm-started with CP or starting from the default ordered list. The last method predecessor-flow follows our outlined flow in Fig. 4). It starts a local search from the default ordered list, and if it does not find a feasible solution within 15 seconds (giving local search sufficient time to initialize), it runs a CP warm start instead.
leftmost leftmost-CP predecessor predecessor-CP predecessor-flow % feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ GEN-1 21.4 583.5 0.0 100.0 802.1 0.0 0.1 651.0 0.0 85.9 793.0 0.0 85.8 792.2 0.0 GEN-0.98 100.0 312.8 0.0 89.4 625.5 0.0 100.0 143.4 3.0 73.3 341.9 0.0 99.0 149.4 3.2 GEN-0.96 100.0 259.6 0.0 98.3 511.4 0.0 100.0 29.4 26.0 83.7 144.6 0.6 100.0 33.7 25.0 GEN-0.94 100.0 232.7 0.0 99.1 437.9 0.0 100.0 6.5 51.8 86.4 65.4 2.4 100.0 7.7 50.6 GEN-0.9 100.0 190.7 0.0 100.0 326.4 0.0 100.0 0.9 80.2 87.4 21.3 13.5 100.0 0.9 81.2 TOP-L 69.1 919.4 0.0 100.0 836.1 0.0 23.7 258.1 0.0 82.8 766.8 0.0 80.8 731.6 0.0 TOP-L-L 100.0 808.2 0.0 85.3 773.8 0.0 100.0 20.2 84.7 75.8 267.3 2.4 99.7 27.8 81.3 TOP-R 69.0 135.6 2.0 99.9 158.4 6.7 64.2 11.0 25.0 98.0 142.0 2.8 99.0 93.6 22.3 TOP-R-L 100.0 101.5 3.9 92.0 113.8 29.3 100.0 0.0 99.7 91.2 8.9 70.9 100.0 0.0 99.8
@@
We can see that the predecessor FFS heuristic achieves great results if the utilization is below 98%. It finds a feasible solution and achieves low values of $D_{sum}$. In case of predecessor-CP, CP solver finds a feasible schedule in most cases (99% for when solving one resource after another), but it may not produce an ordered list acceptable by the FFS heuristic (see Section 4.2.2). Even when the ordered list succeeds, final values of $D_{sum}$ seem to be worse than in the case of predecessor results. The reason may be that the initial ordered list lacks rate-monotonicity, making it harder for local search to improve it. However, CP warm start helps to solve the most utilized instances among our sets TOP-L and GEN-1.
@@
Finally, we also tinkered with using the genetic algorithm to find a suitable ordered list of tasks. However, we did not pursue it further; the ordered lists obtained by the genetic algorithm’s operators too often produced infeasible solutions, and within the same number of evaluations, the genetic algorithm found either worse or no solution at all. Given the constrained characteristics of our problem and the use of the FFS heuristic, the local search appears to be much more efficient.
@@ 5.2 Benefit of Matheuristic
We compared two ways how to use matheuristic: (i) fixed matheuristic, when the matheuristic was initiated after five minutes of local search; and (ii) dynamic matheuristic, when the matheuristic starts once the local search does not improve the solution for 60 seconds. We determined this dynamic threshold from the small experiment, studying how the local search improves the solution over time (see Fig. 7(a)). The local search with no matheuristic is used as a baseline. In all cases, the computation time was six minutes.
In Table 3, we compare the degeneracy obtained while using fixed or dynamic matheuristic ($D_{sum}^{\mathrm{MH}}$) and the degeneracy obtained by the local search alone ($D_{sum}$). We report averaged values of $D_{sum}^{\mathrm{MH}}$ and differences $\Delta D_{sum}=D_{sum}^{\mathrm{MH}}-D_{sum}$. Thus, a negative value of $\Delta D_{sum}$ means that, on average, using matheuristic was beneficial.
fixed matheuristic dynamic matheuristic avg $D_{sum}^{\mathrm{MH}}$ avg $\Delta D_{sum}$ avg $D_{sum}^{\mathrm{MH}}$ avg $\Delta D_{sum}$ GEN-1 642.0 -151.0 687.0 -106.0 GEN-0.98 128.1 -15.3 130.2 -12.6 GEN-0.96 32.3 2.9 27.1 -2.3 GEN-0.94 6.4 -0.1 6.4 -0.1 GEN-0.9 0.9 0.0 1.0 0.1 TOP-L 585.8 -180.2 684.4 -45.2 TOP-L-L 23.8 3.6 25.5 -2.3 TOP-R 55.8 -86.6 63.2 -29.7 TOP-R-L 0.0 0.0 0.0 0.0
@@
Especially for the maximally-utilized instances, the benefit of the matheuristic is significant - the newly obtained values of $D_{sum}^{\mathrm{MH}}$ are significantly lower than the previous values $D_{sum}$. However, for the fixed matheuristic, its benefit drops significantly and even diminishes for some instances - the local search probably has not yet converged, and running it further could yield a better solution. This can be seen in Fig. 9. The histogram of values $\Delta D_{sum}$ is shown for instances GEN-0.96 (for clarity, only those with $\Delta D_{sum}\neq 0$ are presented). We can see that the majority of the differences are negative, meaning the matheuristic was beneficial. However, a few instances with very large positive values of $\Delta D_{sum}$ skew the average to the reported value of $2.9$.
@@
While the dynamic setting does not produce as good results as the fixed one for the most highly-utilized instances, we observed that its results are more consistent across the instance sets; e.g., on GEN-0.98, the worst result reported by the fixed setting was the value of $D_{sum}$ greater by 253 than the one reported by the local search only. In contrast, the worst difference reported by dynamic matheuristic was only 19. Therefore, we used the dynamic matheuristic in the rest of the paper.

### v2

@@ 4.3 Constraint Programming Matheuristic
$\mathtt{EndOf}(J_{i}^{k})\leq\tau(\beta)$ $\forall{m}\in\left\{1,\dots,M\right\},\leavevmode\nobreak\ \forall J_{i}^{k}\in\mathcal{J}_{m}^{O}$ (59)
$ps_{i}^{k}=(\mathtt{StartOf}(J_{i}^{k})+o_{\pi_{i}^{k}})\mod\tau(\beta)$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)},\leavevmode\nobreak\ \forall i\in\left\{1,\dots,n^{k}\right\}$ (60)
$pe_{i}^{k}=(\mathtt{EndOf}(J_{i}^{k})+o_{\pi_{i}^{k}})\mod\tau(\beta)$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)},\leavevmode\nobreak\ \forall i\in\left\{1,\dots,n^{k}\right\}$ (61)
${df}_{0}^{k}=0$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)}$ (62)
$pd_{i+1}^{k}=(ps_{i+1}^{k}-pe_{i}^{k})<0$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)},$
$\forall i\in\left\{1,\dots,n^{k}-1\right\}$ (63)
${df}_{i+1}^{k}=ps_{i+1}^{k}-pe_{i}^{k}+\tau(\beta)\cdot pd_{i+1}^{k}$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)},$
$\forall i\in\left\{1,\dots,n^{k}-1\right\}$ (64)
$S^{k}=\sum_{\forall J_{i}^{k}\in\mathcal{C}^{k}}p_{i}^{k}+{df}_{i}^{k}$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)}$ (65)
$J_{i}^{k}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathtt{intervalVar}$ $\forall{m}\in\left\{1,\dots,M\right\},$
$\forall J_{i}^{k}\in\mathcal{J}_{m}^{F}\cup\mathcal{J}_{m}^{O}$ (66)
$S^{k}\in\mathbb{Z}$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)}$ (67)
$ps_{i}^{k},pe_{i}^{k},df_{i}^{k}\in\mathbb{Z}$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)},\leavevmode\nobreak\ \forall i\in\left\{1,\dots,n^{k}\right\}$ (68)
$pd_{i}^{k}\in\left\{0,1\right\}$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)},\leavevmode\nobreak\ \forall i\in\left\{1,\dots,n^{k}\right\}$ (69)
@@
Eq. 56 ensures the non-overlapping of intervals on each resource. Eqs. 57, 58 and 59 either fix or limit the intervals of the tasks (their occurrences). Variables $ps_{i}^{k}$ and $pe_{i}^{k}$ correspond to the start- and completion time of the task once the resource offset is considered. These (with helper binary variables $pd_{i}^{k}$) are used to calculate $df_{i}^{k}$ the consecutive differences between completion- and start times of the tasks in Eqs. 63 and 64. These, together with tasks’ processing times, form the chain’s start-to-end latency $S^{k}$ in equation Eq. 65. As in the case of the FFS heuristic, Eqs. 63 and 64 regarding the calculation of $df_{i}^{k}$ could be modified to account for link delays or cut-through mechanisms.
@@
The objective function precisely models the degeneracy $D_{sum}$. Note that both $D_{\max}$ and $D_{sum}({\alpha})$ can be modeled in a similar manner. Furthermore, the following constraint Eq. 70 is used with sum-based objectives to ensure that the degeneracy of individual chains is not worsened. Let $D^{k}$ be the degeneracy of chain $\mathcal{C}^{k}$ given the current solution. Then, the newly added constraints are:
$S^{k}\leq(D^{k}+1)\cdot\tau(\beta)$ $\forall\mathcal{C}^{k}\in{C}^{\tau(\beta)}$ (70)
@@
For large instances, the performance of the CP solver starts to diminish due to the large number of task intervals; the solver cannot reduce the degeneracy of any chain even though the search space is large. Thus, when the size of the problem exceeds a defined bound, not all chains are optimized simultaneously; rather, we optimize only $K$ chains with the largest value of degeneracy (keeping others fixed). With this reduction, the solver can again improve the objective within the provided time. The process is repeated (each time with $K$ most problematic chains) until a time limit is reached.
@@ 5 Experiments
The experiments were primarily performed on synthetically generated instances inspired by those in [35]. Initially, each resource is split into segments that become tasks of period $\tau(1)$ and segments that are recursively split further to allocate longer-period tasks. Base period $\tau(1)$ was set to 100, 200, or 400, and there were between 3 to 5 distinct periods per instance, with ratios of 2, 3, and 4 between consecutive periods. This ensures that each instance is feasible. Chains are generated afterward, ensuring that the degeneracy of any synthetic instance is zero. Thus, the synthetic instances generated in this manner all admit a feasible solution with an optimal degeneracy of zero, making the comparison of the methods easy.
@@
Figure 9: Topologies used for TOP (a, b, c) and LINE (d) instance sets.
Three kinds of instance sets were produced. Set GEN-x contains the instances for the general case of PSP with 5 or 10 resources. Chains can pass from one arbitrary resource to another, even returning to the original one; the support graph of such an instance can be a complete graph if the instance’s chains induce it. This corresponds to the most general case of PSP. Furthermore, the processing times of tasks of a single chain can differ. x refers to $\min_{m}U_{m}$; the minimum utilization of resources.
Sets TOP and LINE were generated using four different underlying topologies: star, triangle, bridge, and stubbed line (used topologies are shown in Fig. 9). LINE contains instances using the stubbed line topology; the rest is put into the set TOP. This was done so the results on more complex stubbed-line instances with longer chains would not be skewed by the instances generated using smaller topologies. Processing time is the same for all tasks of each chain. The links are full-duplex, and messages are sent between endpoints. During generation, two random endpoint network nodes were sampled, and a chain was created between them if possible. This chain sampling was repeated until no more chains with zero degeneracy could be scheduled. Furthermore, there are also tasks being sent between two neighboring nodes only (essentially forming a chain of length 1) to increase the utilization. These additional tasks were omitted to obtain less-utilized sets TOP-LESS, LINE-LESS).
Finally, THEORY set is used to discuss the complexity results; it consists of acyclic instances with a bottleneck resource. Specifically, these instances correspond to scattering cases with line and tree support graphs (see Section 3.2). The parameters of the instance sets are presented in Table 1.
# instances # resources avg # tasks avg # chains avg chain length avg $\min_{m}U_{m}$ GEN-1 1194 5-10 2193 254 8.4 1.00 GEN-0.98 1194 5-10 1786 208 8.3 0.98 GEN-0.96 1194 5-10 1558 183 8.2 0.96 GEN-0.94 1194 5-10 1374 163 8.1 0.94 GEN-0.9 1194 5-10 1093 131 7.9 0.90 LINE 796 46 7634 2247 2.7 1.00 LINE-LESS 796 46 6155 2246 2.7 0.71 TOP 2388 10-24 2487 792 2.3 1.00 TOP-LESS 2388 10-24 1929 792 2.35 0.73 THEORY 320 10-77 1102 96 11.2 0.66
@@ 5.1 Finding Feasible Solutions
First, we studied the effect of a warm start and the ability of the local search to find feasible and high-quality solutions. We ended with a strategy that uses the CP to warm-start the heuristic for instances with utilization close to 100%; otherwise, it relies on the local search itself. This strategy provided a good trade-off between the feasibility success rate and the value of the objective function.
The experiment was performed on TOP, LINE, and GEN instances with a time limit of 6 minutes. In Table 2, we report the percentage of instances where the feasible solution was found, the median (and standard deviation) of degeneracy $D_{sum}$, and a percentage of instances with 0-$D_{sum}$ solutions (when the final solution has degeneracy equal to 0). Note that the values of $D_{sum}$ are reported only for a subset of instances for which the method found any solution at all. Thus, it is necessary to compare both the ability to find feasible solutions and the ability to optimize them efficiently when analyzing the results so that a practically inapplicable method that is efficient in a few easy instances does not skew the conclusions. Finally, we also show the median of $D_{sum}^{chains}$, which is the degeneracy $D_{sum}$ divided by the number of chains. This metric describes the delay of a mean chain; however, different lengths of chains and their number may skew the resulting value.
The results are reported for both leftmost and predecessor heuristic methods, either being warm-started with CP or starting from the default ordered list. The last method predecessor-flow follows Fig. 5). It starts a local search from the default ordered list, and if it does not find a feasible solution within 15 seconds (giving local search sufficient time to initialize), it runs a CP warm start instead (Section 4.2).
leftmost leftmost-CP % feasible $D_{sum}$ $D_{sum}^{chains}$ 0-$D_{sum}$ feasible $D_{sum}$ $D_{sum}^{chains}$ 0-$D_{sum}$ GEN-1 21.4 378.5 (538.2) 2.1 (0.6) 0.0 100.0 621.5 (624.3) 3.0 (0.4) 0.0 GEN-0.98 100.0 220.0 (328.4) 1.3 (0.4) 0.0 89.4 467.0 (511.6) 2.6 (0.6) 0.0 GEN-0.96 100.0 198.0 (236.8) 1.3 (0.3) 0.0 98.3 369.5 (430.7) 2.5 (0.6) 0.0 GEN-0.94 100.0 181.5 (204.3) 1.3 (0.3) 0.0 99.1 321.0 (372.8) 2.3 (0.6) 0.0 GEN-0.9 100.0 151.0 (155.8) 1.3 (0.4) 0.0 100.0 236.0 (285.3) 2.1 (0.7) 0.0 LINE 69.1 889.5 (326.0) 0.4 (0.0) 0.0 100.0 775.0 (268.4) 0.4 (0.0) 0.0 LINE-LESS 100.0 721.0 (334.3) 0.4 (0.0) 0.0 85.3 720.0 (272.9) 0.3 (0.0) 0.0 TOP 69.0 75.0 (165.1) 0.1 (0.1) 2.0 99.9 110.5 (166.2) 0.2 (0.1) 6.7 TOP-LESS 100.0 48.0 (128.1) 0.1 (0.1) 3.9 92.0 55.0 (143.8) 0.1 (0.1) 29.3
predecessor predecessor-CP predecessor-flow % feasible $D_{sum}$ $D_{sum}^{chains}$ 0-$D_{sum}$ feasible $D_{sum}$ $D_{sum}^{chains}$ 0-$D_{sum}$ feasible $D_{sum}$ $D_{sum}^{chains}$ 0-$D_{sum}$ GEN-1 0.1 651.0 (0.0) 2.3 (0.0) 0.0 85.9 610.0 (626.8) 3.0 (0.4) 0.0 85.8 611.0 (621.3) 3.0 (0.4) 0.0 GEN-0.98 100.0 90.0 (181.2) 0.6 (0.4) 3.0 73.3 256.0 (286.0) 1.5 (0.5) 0.0 99.0 87.5 (201.7) 0.6 (0.5) 3.2 GEN-0.96 100.0 11.0 (69.5) 0.1 (0.2) 26.0 83.7 106.0 (122.3) 0.8 (0.4) 0.6 100.0 11.0 (86.0) 0.1 (0.2) 25.0 GEN-0.94 100.0 0.0 (11.1) 0.0 (0.1) 51.8 86.4 46.0 (58.8) 0.4 (0.3) 2.4 100.0 0.0 (21.3) 0.0 (0.1) 50.6 GEN-0.9 100.0 0.0 (2.6) 0.0 (0.0) 80.2 87.4 12.0 (28.5) 0.1 (0.2) 13.5 100.0 0.0 (2.6) 0.0 (0.0) 81.2 LINE 23.7 243.0 (152.4) 0.1 (0.0) 0.0 82.8 719.0 (231.0) 0.4 (0.0) 0.0 80.8 684.0 (261.8) 0.4 (0.1) 0.0 LINE-LESS 100.0 0.0 (64.5) 0.0 (0.0) 84.7 75.8 147.0 (284.3) 0.1 (0.1) 2.4 99.7 0.0 (78.2) 0.0 (0.0) 81.3 TOP 64.2 2.0 (27.3) 0.0 (0.0) 25.0 98.0 101.0 (148.1) 0.2 (0.1) 2.8 99.0 9.0 (135.3) 0.0 (0.1) 22.3 TOP-LESS 100.0 0.0 (0.1) 0.0 (0.0) 99.7 91.2 0.0 (35.4) 0.0 (0.0) 70.9 100.0 0.0 (0.0) 0.0 (0.0) 99.8
@@
Firstly, the leftmost-based methods are exceptional at finding feasible solutions, especially when the CP warm start is used. However, since they do not consider chains during scheduling, their degeneracy values are larger, even more so once the utilization of instances decreases. Therefore, they are not suitable in our TSN-like scenario. Notice, however, that the results of leftmost-CP for GEN-1 instance set report almost the same values of degeneracy as the other methods.
We can see that the predecessor FFS heuristic achieves great results if the utilization is below 98%. It finds a feasible solution and achieves low values of $D_{sum}$, as well as $D_{sum}^{chains}$. In the case of predecessor-CP, bin-packing CP solver finds a feasible schedule for a warm start in most cases , but it may not produce an ordered list acceptable by the FFS heuristic (see Section 4.2.2). Even when the ordered list succeeds, final values of $D_{sum}$ seem to be worse than in the case of predecessor results. The reason may be that the initial ordered list lacks rate-monotonicity, making it harder for local search to improve it. However, CP warm start helps to solve the most utilized instances among our sets LINE and GEN-1.
@@ 5.2 Benefit of Matheuristic
We compared two ways how to use matheuristic: (i) fixed matheuristic, when the matheuristic was initiated after five minutes of local search; and (ii) dynamic matheuristic, when the matheuristic starts once the local search does not improve the solution for 60 seconds. We determined this dynamic threshold from the behavior of local search (its convergence) on a subset of instances (see Fig. 8(a)), so the matheuristic would not start prematurely. The local search with no matheuristic is used as a baseline. In all cases, the computation time was six minutes.
In Table 3, we compare the degeneracy obtained while using fixed or dynamic matheuristic ($D_{sum}^{\mathrm{MH}}$) and the degeneracy obtained by the local search alone ($D_{sum}$). We report median (and standard deviation) values of $D_{sum}^{\mathrm{MH}}$ and differences $\Delta D_{sum}=D_{sum}^{\mathrm{MH}}-D_{sum}$. Thus, a negative value of $\Delta D_{sum}$ means that, on average, using matheuristic was beneficial.
fixed matheuristic dynamic matheuristic $D_{sum}^{\mathrm{MH}}$ $\Delta D_{sum}$ $D_{sum}^{\mathrm{MH}}$ $\Delta D_{sum}$ GEN-1 458.5 (542.2) -139.0 (93.7) 546.5 (518.4) -26.0 (169.0) GEN-0.98 73.5 (181.4) -12.0 (35.4) 69.5 (178.1) -9.0 (16.0) GEN-0.96 11.0 (83.8) 0.0 (36.8) 11.5 (68.6) 0.0 (7.7) GEN-0.94 0.0 (14.7) 0.0 (11.3) 1.0 (10.3) 0.0 (1.8) GEN-0.9 0.0 (2.5) 0.0 (0.3) 0.0 (2.8) 0.0 (0.5) LINE 552.0 (237.7) -156.0 (65.3) 642.5 (277.9) -28.0 (59.0) LINE-LESS 0.0 (61.3) 0.0 (17.3) 0.0 (71.6) 0.0 (43.3) TOP 34.0 (78.9) -60.0 (89.9) 7.0 (111.5) -3.0 (54.0) TOP-LESS 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
@@
Especially for the maximally-utilized instances, the benefit of the matheuristic is significant - the newly obtained values of $D_{sum}^{\mathrm{MH}}$ are significantly lower than the previous values $D_{sum}$. However, for the fixed matheuristic, its benefit drops significantly and even diminishes for some instances - the local search probably has not yet converged, and running it further could yield a better solution. This can be seen in Fig. 10. The histogram of values $\Delta D_{sum}$ is shown for instances GEN-0.96 (for clarity, only those with $\Delta D_{sum}\neq 0$ are presented). We can see that the majority of the differences are negative, meaning the matheuristic was beneficial. However, a few instances with very large positive values of $\Delta D_{sum}$ skew the average of the difference values to $2.9$ using matheuristic would be detrimental.
@@
While the dynamic setting does not produce as ground-breaking beneficial examples as the fixed one for the most highly-utilized instances, we observed that its results are more consistent across the instance sets; e.g., on GEN-0.98, the most detrimental effect of dynamic matheuristic was $\Delta D_{sum}=+19.0$. This is a much better overall result than the worst case reported for the fixed matheuristic ($+253$) for the same instance set. Therefore, we used the dynamic matheuristic in the rest of the paper.